OpenStudy (loser66):

A roller coaster makes a run down a track from a vertical distance of 25m. If there is negligible friction and the coaster starts from rest, what is its speed at the bottom of the track? Please

OpenStudy (loser66):

@DemolisionWolf

OpenStudy (anonymous):

so no forces but gravity, so, free-fall from rest. it 'falls' 25m can you see which kinematic equation to use now?

OpenStudy (loser66):

W = mgh?

OpenStudy (loser66):

or K =1/2mv^2?

OpenStudy (anonymous):

awhh, so your exam is on energy, kinematics and momentum and newtons forces I take it?

OpenStudy (anonymous):

we'll want to use the kinematic equations for this one. energy uses 'mass' and we arn't given mass. V = V + at v^2 = v^2 + as s = s + vt + (1/2)at^2 so which of these 3 kinematic equations do you know all the variables Except the one you are looking for [velocity]

OpenStudy (anonymous):

haha, no i'm not in your class ^_^ i'm you professor! lol

OpenStudy (loser66):

hey, why?? W = \(\triangle K\) and W = mgh, where K = 1/2mv^2 we cancel out m, no need to have m, right?

OpenStudy (loser66):

Ok @DemolisionWolf my prof !!! hehehe... you are not lucky at all when you have a student like me.

OpenStudy (anonymous):

haha, I see it now, i'm so use to solving 'free-fall' equations with kinetics, not with energy! yaya, you guys are totally right, energy works! weird as it is, the two equations end up looking identical

OpenStudy (loser66):

I am sorry "prof". hehehe.... I will fire back if you say something nonsense to me.

OpenStudy (anonymous):

haha I gotta take off, good luck on your exam!